joe is 8 years older than kathy. 20 years ago he was 3 times as old as kathy. how old is each now
Answer by stanbon(75887)

joe is 8 years older than kathy. 20 years ago he was 3 times as old as kathy. how old is each now
------------------------------------------
Equations:
J = K + 8
(J-20) = 3(K-20)
-----------------------------
Substitute for "J" and solve for "K":
(K+8 -20) = 3K-60
K-12 = 3K-60
2K = 48
K = 24 yrs (Kathy's age now)
------
Since J = K + 8, J = 24+8 = 32 yrs (Joe's age now)
==================================================
Cheers,
Stan H.